OpenStudy (anonymous):

If the sequence is geometric, what is its common ratio?512, 128, 32, 8,

OpenStudy (anonymous):

Hint: \[8=2^3\] \[32=2^5\] \[128=2^7\]

OpenStudy (anonymous):

@destyboobear How can we use that?

OpenStudy (lgbasallote):

here's another hint: \[\large r = \frac{a_2}{a_1}\] that means second term divided by first term